OpenStudy (anonymous):

is the ln(-x) = -1/x dx?

OpenStudy (anonymous):

\[\int\limits (-1/x) dx = - \ln|x|\]

OpenStudy (anonymous):

\[\int\limits_{-3}^{-1}\frac{ du }{ x \sqrt{\ln \left| -x \right|} }\] for this I get 2sqrt ln3 but the answer is -2sqrt ln 3

OpenStudy (anonymous):

you'll need to break up the ln|-x|

OpenStudy (anonymous):

for my u substitution I had u = ln(-x); du = -1/x; which gives me integral of -du/u

OpenStudy (anonymous):

Should I break up the ln(-x) further?

OpenStudy (anonymous):

so, you have ln|-x|, ln(-x) ln|-x| = ln(-x) for -x > 0 ln|-x| = ln(x) for x <= 0 which case are you in?

OpenStudy (anonymous):

I am going from -3 to -1 which means that I would be in the second case (ln|-x| = ln(x) for x <= 0) which means that ln(-x) = ln(x). Is this rigtht?

OpenStudy (anonymous):

it means ln|-x| becomes ln(x).

OpenStudy (anonymous):

I am slightly confused

OpenStudy (anonymous):

so you have: \[\int\limits \frac{dx}{x \ln(x)}\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[\int\limits\limits \frac{dx}{x \sqrt{\ln(x)}}\]

OpenStudy (anonymous):

u = ln(x) du = (1/x) dx, so you have: \[\int\limits \frac{du}{\sqrt{u}}\]

OpenStudy (anonymous):

Did you take out the negative from both the integral and the ln to do that?

OpenStudy (anonymous):

oh crap, I was dead wrong O.O Let's back up ln|-x| = ln(-x) for -x > 0 ln|-x| = ln(x) for -x <= 0 so we're in the first case because -x > 0, means x < 0 so, \[\int\limits \frac{dx}{x \sqrt{\ln(-x)}}\]

OpenStudy (anonymous):

which means that I have u= ln|-x| du = 1/x and integral of du/u. Is this correct?

OpenStudy (anonymous):

Basically what you are doing above is just splitting up the absolute value?

OpenStudy (anonymous):

yes. But you're supposed to let u = ln(-x), du = (1/x)dx the integral becomes 

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I got it now

OpenStudy (anonymous):

which becomes 2 sqrt(ln(-x))

OpenStudy (anonymous):

so you have 2 sqrt(ln(- -1)) - 2 sqrt(ln(- - 3)) = 2 [sqrt(ln(1)) - sqrt(ln(3)) = -2sqrt(ln(3))